# The deep multi-FBSDE method: a robust deep learning method for coupled FBSDEs

### v1

@@ 3.2 The deep multi-FBSDE method
We demonstrate the new numerical scheme on the LQG-problem, detailed in Section 4.1.2, for which the deep BSDE method was shown to fail in Section 3.1. In Phase I, we take $K=1,2,3,4$, with the choices $\psi_{1}=0$, $(\psi_{1},\psi_{2})=(0,b_{1})$, $(\psi_{1},\psi_{2},\psi_{3})=(0,b_{1},-b_{1})$ and $(\psi_{1},\psi_{2},\psi_{3},\psi_{4})=(0,b_{1},-b_{1},-0.5b_{1})$, where the drift has the form $b(x,y,z)=b_{1}(x)+b_{2}(z)$, and plot the MSE curves in Figure 3. For $K=2,3,4$ the optimization landscape is much more satisfactory than for $K=1$, which is the deep BSDE method. In fact the minima are close to the real $Y_{0}=8.94$. To further indicate that the method is robust to the choice of $K\geq 3$ (and possibly $K=2$) and $(\psi_{1},\dots,\psi_{K})$ we take $K=3$ and rather arbitrary choices $\psi_{1}=b_{1}\circ\xi$, $\psi_{2}=-b_{1}\circ\xi$, $\psi_{3}=0$, for $\xi\in\{x,-\sin(x),x\cos(x),1-\exp((x\wedge 1)\vee(-1))\}$, where $\xi$ acts coordinate-wise on vectors. We obtain the four loss surfaces in Figure 4, which, based on visual inspection, have the same minima. This robustness is a very important property because when the method is applied to a problem without a reference solution, one must be able to trust that the choice of $\psi_{1},\psi_{2},\psi_{3}$ does not need to be tuned based on comparison with a reference solution. Our results indicate a robustness in the choice of $K\geq 2$ and $(\psi_{1},\dots,\psi_{K})$, but it needs to be investigated further, particularly theoretically. The deviation between $Y_{0}$ and the optimized $y_{0}^{*}$ is the discretization error due to the time stepping. This is seen empirically in the convergence plot in Figure 5 for $\xi(x)=x$, showing the experimental convergence order one, i.e., $|Y_{0}-y_{0}^{*}(h)|=\mathcal{O}(h)$. The convergence of $(X,Y,Z)$ as processes after Phase II is shown in Section 4.1.2.
@@
Figure 3: Four MSE curves plotted for $K=1,2,3,4$, with $\psi_{1}=0$, $(\psi_{1},\psi_{2})=(0,b_{1})$, $(\psi_{1},\psi_{2},\psi_{3})=(0,b_{1},-b_{1})$ and $(\psi_{1},\psi_{2},\psi_{3},\psi_{4})=(0,b_{1},-b_{1},-0.5b_{1})$, where $b(x,y,z)=b_{1}(x)+b_{2}(z)$. $y_{0}^{*}$ for $K=1$ and $K=2$ are the initial conditions approximated with Phase I or our method. Note that for $K=1$, $y_{0}^{*}=28.44$ as can be visualized in Figure 1. For visualization purposes, we have scaled the curves. For $K=1,2,3,4$, the training times are approximately 400s, 500s, 600s, and 700s, respectively.
@@
Figure 4: Four MSE curves plotted for $\psi_{1}=b\circ\xi$, $\psi_{2}=-b\circ\xi$, $\psi_{3}=0$, for $\xi\in\{x,-\sin(x),x\cos(x),1-\exp((x\wedge 1)\vee(-1))\}$, where $\xi$ acts coordinate wise on vectors. The order coincides with that of the set.
@@ 4 Numerical experiments
We present four numerical experiments, two with an associated stochastic optimal control problem, and two without such a connection, all with analytic or semi-analytic reference solutions so that we can analyze the error. The first problem is a controlled Brownian motion in two dimensions, the second problem is a six-dimensional LQG-problem, and the last two examples are FBSDEs related to two-dimensional advection diffusion reaction PDEs, one without a reaction term and one with a linear one. We demonstrate the performance of both the deep BSDE and the deep multi-FBSDE methods.
@@
In the implementation of the deep BSDE and both Phase I–II of the deep multi-FBSDE methods, we optimize $\zeta_{0},\dots,\zeta_{N-1}$ in a function class defined by ReLU feed-forward neural networks with 3 hidden layers with 20 nodes in each layer. The initial value $y_{0}$ in the deep BSDE method and in Phase I of the deep multi-FBSDE method is represented by one single trainable parameter. Training data consist of $2^{20}$ independent realizations of the $d$-dimensional Wiener increment vectors $\Delta W_{0},\ldots,\Delta W_{N-1}$, each element being i.i.d. and normally distributed with mean zero and variance $h$. The training data are reused in $10$ epochs. Training is initiated by randomly sampling parameters and proceeds with the classical Adam optimizer with batches of size $2^{12}$.
@@ 4.1 FBSDEs stemming from stochastic optimal control problems
In this section, we look into two main types of stochastic control problems, controlled Brownian motions and Linear Quadratic Gaussian (LQG) control problems. Before discussing these specific topics, we define the general form of the stochastic control problem:
@@
Assuming existence, the value function associated with the control problem is formally defined as
@@
Moreover, the value function satisfies the following Hamilton–Jacobi–Bellman (HJB) equation
$\begin{dcases}\frac{\partial v}{\partial t}+\frac{1}{2}\mathrm{Tr}\big{\{}\sigma^{\top}\sigma\mathrm{Hess}_{x}v\big{\}}+\inf_{u}\big{[}b_{u}^{\top}\mathrm{D}_{x}v+f_{u}\big{]}=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ v(t,x)=g(x),&(t,x)\in{T}\times\mathbb{R}^{d}.\end{dcases}$ (13)
The resulting HJB-equation is a semilinear parabolic PDE of the type discussed in Section 2; therefore, it admits an equivalent formulation as an FBSDE.
@@ 4.1.1 Controlled Brownian motions with general terminal cost
$b_{u}(t,x,u)=u,\quad\sigma(t,x)=\sigma I_{d},\quad f(t,x,u)=\frac{r}{2}|u|^{2},$ (14)
@@
$\begin{dcases}X_{t}=x_{0}-\int_{0}^{t}\frac{1}{r}Z_{t}\text{d}s+\sigma W_{t},\\ Y_{t}=g(X_{T})+\int_{t}^{T}\frac{1}{2r}|Z_{s}|^{2}\text{d}s-\int_{t}^{T}\sigma Z_{s}\text{d}W_{s},\quad t\in[0,T].\end{dcases}$ (15)
@@
$Y_{t}=-\frac{1}{r\sigma^{2}}\log\bigg{\{}\mathbb{E}\big{[}\text{e}^{-r\sigma^{2}g(X_{t}+\xi\sqrt{T-t})}\big{|}X_{t}\big{]}\bigg{\}},$ (16)
@@
$\begin{dcases}X_{t}^{\psi_{2}}=x_{0}+\sigma W_{t},\\ Y_{t}^{\psi_{2}}=g(X_{T}^{\psi_{2}})-\int_{t}^{T}\frac{1}{2r}|Z_{s}^{\psi_{2}}|^{2}\text{d}s-\int_{t}^{T}\sigma Z_{s}^{\psi_{2}}\text{d}W_{s},\quad t\in[0,T].\end{dcases}$ (17)
@@
We set $T=0.5$, $d=\ell=k=2$, $r=1$, and $\sigma=0.25$. The initial state is set at $x_{0}=(-0.1,0.1)^{\top}$. We employ a terminal cost function $g(x)=-|x_{1}-x_{2}|$. This configuration aims to control the two components of the Brownian motion such that they diverge as much as possible by the time $t=T$, while minimizing the use of control force. The discretization of the problem is set with $N=40$.
@@
Figure 6 illustrates the loss and the approximation of $Y_{0}$ during training for both the deep FBSDE method and the deep multi-FBSDE method. Notably, while the loss approaches zero for both methods, only the deep multi-FBSDE method acurately converges to the true $Y_{0}$.
@@ 4.1.2 Linear Quadratic Gaussian control problems
Among all stochastic control problems, the linear quadratic Gaussian (LQG) control problem is notably the most structured and extensively studied, see, for example, [4]. For our purposes, it presents a closed-form analytic solution, allowing us to benchmark our numerical approximations against it.
@@
$b_{u}(t,x,u)=A(C-x)+Bu,\quad f_{u}(t,x,u)=\langle R_{x}x,x\rangle+\langle R_{u}u,u\rangle\quad g(x)=\langle Gx,x\rangle.$
The above, together with (12), defines the state equation and cost functional for a LQG control problem. The optimal feedback control, minimizing the Hamiltonian, $\inf_{u\in U}{\langle\text{D}xV,Bu\rangle+\langle R_{u}u,u\rangle}$, is given by
@@
$\begin{dcases}\dot{P}(t)-A^{T}P(t)-P(t)A-P(t)BR_{u}^{-1}B^{T}P(t)+R_{x}=\mathbf{0}_{d\times d},\\ \dot{Q}(t)+2P(t)AC-A^{T}Q(t)-P(t)BR_{u}^{-1}B^{T}Q(t)=\mathbf{0}_{d},\\ \dot{R}(t)+\text{Tr}\big{\{}\sigma\sigma^{\top}P(t)\big{\}}+Q(t)^{\top}AC-\frac{1}{4}Q(t)^{\top}BR_{u}^{-1}B^{\top}Q(t)=0,\quad t\in[0,T],\\ P(T)=G;\quad Q(T)=\mathbf{0}_{d};\quad R(T)=0.\end{dcases}$
@@
$\begin{dcases}X_{t}=x_{0}+\int_{0}^{t}\big{[}A(C-X_{s})-\frac{1}{2}BR_{u}^{-1}B^{\top}Z_{s}\big{]}\text{d}s+\int_{0}^{t}\sigma\text{d}W_{s},\\ Y_{t}=\langle GX_{T},X_{T}\rangle+\int_{t}^{T}\big{(}\langle R_{x}X_{s},X_{s}\rangle+\frac{1}{4}\langle R_{u}^{-1}B^{\top}Z_{s},B^{\top}Z_{s}\rangle\big{)}\text{d}s-\int_{t}^{T}\langle Z_{s},\sigma\text{d}W_{s}\rangle,\quad t\in[0,T].\end{dcases}$ (20)
@@
Figure 8 clearly demonstrates that our method yields accurate approximations for $(X,Y,Z)$, both pathwise and in distribution. We refrain from comparing these results with those obtained via the deep BSDE method, as the latter produce completely incorrect approximations for this problem. The approximate initial condition obtained by the deep BSDE method is around 28.44, whereas the true value is approximately 9 (see Figure 1). In [30], the authors use the deep BSDE method to approximate the solution of a BSDE derived from the stochastic maximum principle for the specific problem discussed above. However, this problem is significantly less complex because, at each time $t$, the control process in the associated BSDE is constant in the state space. Consequently, it remains unclear to what extent this approach would generalize to more complex problems where the control process is non-constant.
@@
Figure 8: Average of solutions and a single solution path compared to their semi-analytic counterparts for the LQG control problem from Section 4.1.2. The shaded area represents an empirical credible interval for $Y$, defined as the area between the 5:th and the 95:th percentiles at each time point. We do not include credible intervals for $X$ and $Z$ in this figure to facilitate visualization. For $X$ and $Z$, we see one realization of each of the six components.
@@ 4.2 FBSDEs from advection diffusion reaction PDEs
Let $a,b\in\mathbb{R}^{+}$, and $x_{0}\in\mathbb{R}^{d}$. For a general function $g\colon\mathbb{R}^{d}\to\mathbb{R}$ we define
$\begin{dcases}\frac{\partial v}{\partial t}(t,x)+a\Delta v+2b|\mathrm{D}_{x}v|^{2}-cv=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ v(t,x)=g(x),&(t,x)\in{T}\times\mathbb{R}^{d}.\end{dcases}$ (22)
The above PDE is structurally similar to a Hamilton–Jacobi–Bellman (HJB) equation. However, since $b>0$, it cannot be written in the form (13), does not have an associate stochastic control problem, and hence does not represent an HJB equation. We consider the equivalent FBSDE
$\begin{dcases}X_{t}=x_{0}+\int_{0}^{t}bZ_{t}\text{d}s+\sqrt{2a}W_{t},\\ Y_{t}=g(X_{T})+\int_{t}^{T}\big{(}b|Z_{s}|^{2}-cY_{s}\big{)}\text{d}s-\int_{t}^{T}\sqrt{2a}Z_{s}\text{d}W_{s},\quad t\in[0,T].\end{dcases}$ (23)
We again employ the Cole–Hopf transformation to derive a reference solution. Specifically, defining $u(t,x)=\log\Big{\{}\frac{a}{2b}v(t,x)\Big{\}}$ leads directly to the simplified PDE
$\begin{dcases}\frac{\partial u}{\partial t}(t,x)+a\Delta u-cu\log{u}=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ u(t,x)=\exp\Big{\{}\frac{2b}{a}g(x)\Big{\}},&(t,x)\in{T}\times\mathbb{R}^{d}.\end{dcases}$ (24)
Back transformation yields $u(t,x)=\exp\Big{\{}\frac{2b}{a}v(t,x)\Big{\}}$. In the special case, $c=0$ we have for $t\in[0,T]$
$Y_{t}=-\frac{2b}{a}\log\bigg{\{}\mathbb{E}\big{[}\text{e}^{-\frac{a}{2b}g(X_{t}+\xi\sqrt{T-t})}\big{|}X_{t}\big{]}\bigg{\}},$ (25)
@@
For the deep multi-FBSDE method, we opt for $K=2$, setting $\Psi_{1}(t,x,y,z)=0$ and $\Psi_{2}(t,x,y,z)=-\frac{b}{2}z$. We employ a terminal condition $g(x)=-|x_{1}-x_{2}|$. The discretization of the problem is set with $N=40$. In our first experiment, we set $a=0.0315$, $b=0.6$, $x_{0}=(-0.1,0.1)^{\top}$, and $c=0$ and use (25) to compute a reference solution. In our second experiment, we set $a=0.0315$, $x_{0}=(-0.1,0.1)^{\top}$, $c=1$ and let $b$ vary between 0.00125 and 5.
For the first experiment, the reference solution is computed with a Monte-Carlo approximation of (25). For the second experiment, we use a finite difference scheme to approximate (24). More specifically, we approximate the solutions on a spatial domain $[-2,2]\times[-2,-2]$ with artificial Dirichlet boundary conditions.
@@
The figure clearly illustrates that the deep BSDE method yields inaccurate approximations, whereas the deep multi-FBSDE method produces highly accurate results. For the second experiment, we explore the accuracy of the deep BSDE method and the deep multi-FBSDE method for different values of $b$. In Figure 11, we see that for $b>0.25$, the deep BSDE fails to provide accurate approximations. The deep multi-FBSDE method provides accurate approximations for all $b$ tested.
Figure 11: Approximate initial values for different values of $b$. To no distrort visibility of the plot we have used the maximum of -1 and the approximate initial conditions ($Y_{0}$ approximated with the deep BSDE method for $b=10$, $b=1.25$ and $b=1$ is $-1.04$, $-2.9$ and $-4.9$ respectively).

### v2

@@ 3.2 The deep multi-FBSDE method
We demonstrate the new numerical scheme on the LQG-problem, detailed in Section 4.1.2, for which the deep BSDE method was shown to fail in Section 3.1. In Phase I, we take $K=1,2,3,4$, with the choices $\psi_{1}=0$, $(\psi_{1},\psi_{2})=(0,b_{1})$, $(\psi_{1},\psi_{2},\psi_{3})=(0,b_{1},-b_{1})$ and $(\psi_{1},\psi_{2},\psi_{3},\psi_{4})=(0,b_{1},-b_{1},-0.5b_{1})$, where the drift has the form $b(x,y,z)=b_{1}(x)+b_{2}(z)$, and plot the MSE curves in Figure 3. For $K=2,3,4$ the optimization landscape is much more satisfactory than for $K=1$, which is the deep BSDE method. In fact the minima are close to the real $Y_{0}=8.94$. To further indicate that the method is robust to the choice of $K\geq 3$ (and possibly $K=2$) and $(\psi_{1},\dots,\psi_{K})$ we take $K=3$ and rather arbitrary choices $\psi_{1}=b_{1}\circ\xi$, $\psi_{2}=-b_{1}\circ\xi$, $\psi_{3}=0$, for $\xi\in\{x,-\sin(x),x\cos(x),1-\exp(\min(1,\max(-1,x)))\}$, where $\xi$ acts coordinate-wise on vectors. We obtain the four loss surfaces in Figure 4, which, based on visual inspection, have the same minima. This robustness is a very important property because when the method is applied to a problem without a reference solution, one must be able to trust that the choice of $\psi_{1},\psi_{2},\psi_{3}$ does not need to be tuned based on comparison with a reference solution. Our results indicate a robustness in the choice of $K\geq 2$ and $(\psi_{1},\dots,\psi_{K})$, but it needs to be investigated further, particularly theoretically. The deviation between $Y_{0}$ and the optimized $y_{0}^{*}$ is the discretization error due to the time stepping. This is seen empirically in the convergence plot in Figure 5 for $\xi(x)=x$, showing the experimental convergence order one, i.e., $|Y_{0}-y_{0}^{*}(h)|=\mathcal{O}(h)$. The convergence of $(X,Y,Z)$ as processes after Phase II is shown in Section 4.1.2.
@@
Figure 3: Four MSE curves plotted for $K=1,2,3,4$, with $\psi_{1}=0$, $(\psi_{1},\psi_{2})=(0,b_{1})$, $(\psi_{1},\psi_{2},\psi_{3})=(0,b_{1},-b_{1})$ and $(\psi_{1},\psi_{2},\psi_{3},\psi_{4})=(0,b_{1},-b_{1},-0.5b_{1})$, where $b(x,y,z)=b_{1}(x)+b_{2}(z)$. In Phase II $y_{0}^{*}$ for $K=1$ and $K=2$ are the initial conditions approximated with Phase I. Note that for $K=1$, $y_{0}^{*}=28.44$ as can be visualized in Figure 1. For visualization purposes, we have scaled the curves. For $K=1,2,3,4$, the training times are approximately 400s, 500s, 600s, and 700s, respectively.
@@
Figure 4: Four MSE curves plotted for $\psi_{1}=b_{1}\circ\xi$, $\psi_{2}=-b_{1}\circ\xi$, $\psi_{3}=0$, for $\xi\in\{x,-\sin(x),x\cos(x),1-\exp((x\wedge 1)\vee(-1))\}$, where $\xi$ acts coordinate wise on vectors. The order coincides with that of the set.
@@ 4 Numerical experiments
We present four numerical experiments, two with an associated stochastic optimal control problem, and two without such a connection, all with analytic or semi-analytic reference solutions so that we can analyze the error. The first problem is a controlled Brownian motion in two dimensions, the second problem is a six-dimensional LQ-problem, and the last two examples are FBSDEs related to two-dimensional advection diffusion reaction PDEs, one without a reaction term and one with a linear one. We demonstrate the performance of both the deep BSDE and the deep multi-FBSDE methods.
@@
In the implementation of the deep BSDE and both Phase I–II of the deep multi-FBSDE methods, we optimize $\zeta_{0},\dots,\zeta_{N-1}$ in a function class defined by ReLU feed-forward neural networks with 3 hidden layers with 20 nodes in each layer. The initial value $y_{0}$ in the deep BSDE method and in Phase I of the deep multi-FBSDE method is represented by one single trainable parameter. Training data consist of $2^{20}$ independent realizations of the $d$-dimensional Wiener increment vectors $\Delta W_{0},\ldots,\Delta W_{N-1}$, each element being i.i.d., and normally distributed with mean zero and variance $h$. The training data are reused in $10$ epochs. Training is initiated by randomly sampling parameters and proceeds with the classical Adam optimizer with batches of size $2^{12}$.
@@ 4.1 FBSDEs stemming from stochastic optimal control problems
In this section, we look into two main types of stochastic control problems, controlled Brownian motions and linear quadratic Gaussian control problems. Before discussing these specific topics, we define the general form of the stochastic control problem:
@@
The value function associated with the control problem is formally defined as
@@
Under sufficient regularity assumptions on $b,\sigma,f,g$, a solution to the following Hamilton–Jacobi–Bellman (HJB) equation exists and coincides with the value function of the control problem
$\begin{dcases}\frac{\partial v}{\partial t}+\frac{1}{2}\mathrm{Tr}\big{(}\sigma^{\top}\sigma\mathrm{Hess}_{x}v\big{)}+\inf_{u}\big{[}\langle\mathrm{D}_{x}v,b\rangle+f\big{]}=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ v(T,x)=g(x),&x\in\mathbb{R}^{d}.\end{dcases}$ (13)
In this paper we tacitly assume such regularity. The resulting HJB-equation is a semilinear parabolic PDE of the type discussed in Section 2; therefore, it admits an equivalent formulation as an FBSDE.
@@ 4.1.1 Controlled Brownian motions with general terminal cost
$b(t,x,u)=u,\quad\sigma(t,x)=\sigma I_{d},\quad f(t,x,u)=\frac{r}{2}\|u\|^{2},$ (14)
@@
$\begin{dcases}X_{t}=x_{0}-\int_{0}^{t}\frac{1}{r}Z_{s}\text{d}s+\sigma W_{t},\\ Y_{t}=g(X_{T})+\int_{t}^{T}\frac{1}{2r}\|Z_{s}\|^{2}\text{d}s-\int_{t}^{T}\langle Z_{s},\sigma\text{d}W_{s}\rangle,\quad t\in[0,T].\end{dcases}$ (15)
@@
$Y_{t}=-\frac{1}{r\sigma^{2}}\log\bigg{(}E\big{[}\text{e}^{-r\sigma^{2}g(X_{t}+\xi\sqrt{T-t})}\big{|}X_{t}\big{]}\bigg{)},$ (16)
@@
$\begin{dcases}X_{t}^{\psi_{2}}=x_{0}+\sigma W_{t},\\ Y_{t}^{\psi_{2}}=g(X_{T}^{\psi_{2}})-\int_{t}^{T}\frac{1}{2r}\|Z_{s}^{\psi_{2}}\|^{2}\text{d}s-\int_{t}^{T}\big{\langle}Z_{s}^{\psi_{2}},\sigma\text{d}W_{s}\big{\rangle},\quad t\in[0,T].\end{dcases}$ (17)
@@
We set $T=0.5$, $d=\ell=k=2$, $r=1$, and $\sigma=0.25$. The initial state is set at $x_{0}=(-0.1,0.1)^{\top}$. We employ a terminal cost function $g(x)=-|x_{1}-x_{2}|$, where $x=(x_{1},x_{2})^{\top}$. This configuration aims to control the two components of the Brownian motion such that they diverge as much as possible by the time $t=T$, while minimizing the use of control force. The discretization of the problem is set with $N=40$.
@@
Figure 6 illustrates the loss and the approximation of $Y_{0}$ during training for both the deep FBSDE method and the deep multi-FBSDE method. Notably, while the loss approaches zero for both methods, only the deep multi-FBSDE method converges to the true $Y_{0}$.
@@ 4.1.2 Linear Quadratic Gaussian control problems
Among all stochastic control problems, the linear quadratic Gaussian control problem is notably the most structured and extensively studied, see, for example, [4]. For our purposes, it presents a closed-form analytic solution, allowing us to benchmark our numerical approximations against it.
@@
$b(t,x,u)=A(C-x)+Bu,\quad f(t,x,u)=\langle R_{x}x,x\rangle+\langle R_{u}u,u\rangle\quad g(x)=\langle Gx,x\rangle.$
The above, together with (12), defines the state equation and cost functional for a LQ control problem. The optimal feedback control, minimizing the Hamiltonian, $\inf_{u\in U}{\langle\text{D}xV,Bu\rangle+\langle R_{u}u,u\rangle}$, is given by
@@
$\begin{dcases}\dot{P}(t)-A^{T}P(t)-P(t)A-P(t)BR_{u}^{-1}B^{T}P(t)+R_{x}=\mathbf{0}_{d\times d},\\ \dot{Q}(t)+2P(t)AC-A^{T}Q(t)-P(t)BR_{u}^{-1}B^{T}Q(t)=\mathbf{0}_{d},\\ \dot{R}(t)+\text{Tr}\big{(}\sigma\sigma^{\top}P(t)\big{)}+Q(t)^{\top}AC-\frac{1}{4}Q(t)^{\top}BR_{u}^{-1}B^{\top}Q(t)=0,\quad t\in[0,T],\\ P(T)=G;\quad Q(T)=\mathbf{0}_{d};\quad R(T)=0.\end{dcases}$
@@
$\begin{dcases}X_{t}=x_{0}+\int_{0}^{t}\big{[}A(C-X_{s})-\frac{1}{2}BR_{u}^{-1}B^{\top}Z_{s}\big{]}\text{d}s+\int_{0}^{t}\sigma\text{d}W_{s},\\ Y_{t}=\langle GX_{T},X_{T}\rangle+\int_{t}^{T}\big{(}\langle R_{x}X_{s},X_{s}\rangle+\frac{1}{4}\big{\langle}R_{u}^{-1}B^{\top}Z_{s},B^{\top}Z_{s}\big{\rangle}\big{)}\text{d}s-\int_{t}^{T}\langle Z_{s},\sigma\text{d}W_{s}\rangle,\quad t\in[0,T].\end{dcases}$ (20)
@@
Figure 8 clearly demonstrates that our method yields accurate approximations for $(X,Y,Z)$, both pathwise and in distribution. We refrain from comparing these results with those obtained via the deep BSDE method, as the latter produces completely incorrect approximations for this problem. The approximate initial condition obtained by the deep BSDE method is around 28.44, whereas the true value is approximately 9 (see Figure 1). In [30], the authors use the deep BSDE method to approximate the solution of a BSDE derived from the stochastic maximum principle for the specific problem discussed above. However, this problem is significantly less complex because, at each time $t$, the control process in the associated BSDE is constant in the state space. Consequently, it remains unclear to what extent this approach would generalize to more complex problems where the control process is non-constant.
@@
Figure 8: Average of solutions and a single solution path compared to their semi-analytic counterparts for the LQ-control problem from Section 4.1.2. The shaded area represents an empirical credible interval for $Y$, defined as the area between the 5:th and the 95:th percentiles at each time point. We do not include credible intervals for $X$ and $Z$ in this figure to facilitate visualization. For $X$ and $Z$, we see one realization of each of the six components.
@@ 4.2 FBSDEs from advection diffusion reaction PDEs
Let $\alpha,\beta\in\mathbb{R}^{+}$, $\gamma\in\mathbb{R}$, and $x_{0}\in\mathbb{R}^{d}$. For a general function $g\colon\mathbb{R}^{d}\to\mathbb{R}$ we define
$\begin{dcases}\frac{\partial v}{\partial t}+\alpha\Delta v+2\beta\|\mathrm{D}_{x}v\|^{2}-\gamma v=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ v(T,x)=g(x),&x\in\mathbb{R}^{d}.\end{dcases}$ (22)
The above PDE is structurally similar to a Hamilton–Jacobi–Bellman (HJB) equation. However, since $\beta>0$, it cannot be written in the form (13), does not have an associate stochastic control problem, and hence does not represent an HJB equation. We consider the equivalent FBSDE
$\begin{dcases}X_{t}=x_{0}+\int_{0}^{t}\beta Z_{t}\text{d}s+\sqrt{2\alpha}W_{t},\\ Y_{t}=g(X_{T})+\int_{t}^{T}\big{(}\beta\|Z_{s}\|^{2}-\gamma Y_{s}\big{)}\text{d}s-\sqrt{2a}\int_{t}^{T}\langle Z_{s},\text{d}W_{s}\rangle,\quad t\in[0,T].\end{dcases}$ (23)
We again employ the Cole–Hopf transformation to derive a reference solution. Specifically, defining $u(t,x)=\log\big{(}\frac{\alpha}{2\beta}v(t,x)\big{)}$ leads directly to the simplified PDE
$\begin{dcases}\frac{\partial u}{\partial t}(t,x)+\alpha\Delta u-cu\log{u}=0,&(t,x)\in[0,T)\times\mathbb{R}^{d},\\ u(T,x)=\exp\Bigg{(}\frac{2\beta}{\alpha}g(x)\Bigg{)},&x\in\mathbb{R}^{d}.\end{dcases}$ (24)
Back transformation yields $u(t,x)=\exp(\frac{2\beta}{\alpha}v(t,x))$. In the special case $\gamma=0$, we have for $t\in[0,T]$
$Y_{t}=-\frac{2\beta}{\alpha}\log{\bigg{(}E\Big{[}\text{e}^{-\frac{\alpha}{2\beta}g(X_{t}+\xi\sqrt{T-t})}\big{|}X_{t}\Big{]}\bigg{)}},$ (25)
@@
For the deep multi-FBSDE method, we opt for $K=2$, setting $\psi_{1}(t,x,y,z)=0$ and $\psi_{2}(t,x,y,z)=-\frac{\beta}{2}z$. We employ a terminal condition $g(x)=-|x_{1}-x_{2}|$, where $x=(x_{1},x_{2})^{\top}$. The discretization of the problem is set with $N=40$. In our first experiment, we set $\alpha=0.0315$, $\beta=0.6$, $x_{0}=(-0.1,0.1)^{\top}$, and $\gamma=0$ and use (25) to compute a reference solution. In our second experiment, we set $\alpha=0.0315$, $x_{0}=(-0.1,0.1)^{\top}$, $\gamma=1$, and let $\beta$ vary between 0.00125 and 5.
For the first experiment, the reference solution is computed with a Monte Carlo approximation of (25). For the second experiment, we use a finite difference scheme to approximate (24). More specifically, we approximate the solutions on a spatial domain $[-2,2]\times[-2,2]$ with artificial Dirichlet boundary conditions.
@@
The figure clearly illustrates that the deep BSDE method yields inaccurate approximations, whereas the deep multi-FBSDE method produces highly accurate results. For the second experiment, we explore the accuracy of the deep BSDE method and the deep multi-FBSDE method for different values of $\beta$. In Figure 11, we see that for $\beta>0.25$, the deep BSDE fails to provide accurate approximations. The deep multi-FBSDE method provides accurate approximations for all $\beta$ tested.
Figure 11: Approximate initial values for different values of $\beta$. To not distort visibility of the plot we have used the maximum of $-1$ and the approximate initial conditions ($Y_{0}$ approximated with the deep BSDE method for $\beta=10$, $\beta=1.25$ and $\beta=1$ is $-1.04$, $-2.9$ and $-4.9$, respectively).